OpenStudy (anonymous):

Write the slope-intercept equation for the line that passes through (-4, -8) and (-3, -2) is this correct? y2-y1/x2-x1 -2-(-8)/-3-(-4)=6/1

OpenStudy (anonymous):

yes thats correct

OpenStudy (anonymous):

do i need to write in y=mx+b as my final answer

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

so yu will do y=6x+b and then solve for b

OpenStudy (anonymous):

yu know how to, right?

OpenStudy (anonymous):

yes i was just about to ask you did i have to solve for b :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

use one of the given points to solve for b

OpenStudy (anonymous):

16=b

OpenStudy (anonymous):

say (-4,-8) -8=6(-4)+b b=16

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

so, y=6x+16